What is the maximum distance we can shoot a dart, from ground level, provided our toy dart gun gives a maximum initial velocity of 3.55 m/s and air resistance is negligible

A. 12.6 m

B. 2.57 m
C. 0.642 m
D. 1.28 m
Answer #1

maximum Range= V^2/g
         =3.55^2/9.8
         =1.28 m
Ans:D
Answer #3

d = v2 sin(2*theta)/g

it will travel the maximum distance for theta = 45 degree

--> d max = v2 sin(2*45)/g

= (3.55^2)/9.8

= 1.285 m